[figure: 9.7 cm]

VOLUME EFFECTS ON ELECTRONIC TRANSITIONS IN Pr_{0.7} Ca_{0.3} MnO_{3} – BASED PEROVSKITES

I.Medvedeva^{a}, A.Maignan^{b}, K.Bärner^{c}, B.Raveau^{b} aInstituite for Metal Physics, Ural Division of RAS, S.Kovalevskaya 18,620219 Ekaterinburg, Russia

bLaboratoire CRISTMAT, UMR 6508 associee au CNRS, Bulevard du Marechal Juin,

14050 Caen, France

cIV Physikalisches Institut, Georg–August–Universität Göttingen, Bunsenstrasse 11-15, D-37073 Göttingen, Germany

The volume effect on the metal-insulator transition temperature T_{mi} and the charge ordering transition temperature T_{co} in Pr_{0.7}Ca_{0.3}MnO_{3}–based compounds was investigated. Different kinds of lattice modulation were used – the average cation size was changed via La substitution for Pr and Sr substitution for Ca [1]; hydrostatic pressure was applied to the resulting Pr_{0.7−x} La_{x} Ca_{0.3}MnO_{3} and Pr_{0.7} Ca_{0.3−x}Sr_{x}MnO_{3} systems. From the temperature dependences of the resistivity the pressure derivatives dT_{mi}/dP and dT_{co}/dP were determined. The main result of this study is evidence for a critical state of Pr_{0.66} Ca_{0.34} MnO_{3} –based systems which is extremely sensitive to external forces, such as hydrostatic pressure, magnetic field and A-cation substitution (chemical pressure).

[1] A.Maignan, Ch.Simon, V.Caignaert, B.Raveau, Z.Phys.B.99 (1996) 305

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : K.Bärner

Address for correspondence : IV Physikalisches Institut, Georg–August–Universität Göttingen, Bunsenstrasse 11-15, D-37073 Göttingen, Germany

Email address : baerner@physik4.gwdg.de